OpenStudy (anonymous):

Determine which angle is such that it's supplement is four times the angle itself. * I'm so lost on this, please help. :\

OpenStudy (anonymous):

x=4(180-x) solve for x

OpenStudy (anonymous):

Let x be the angle. x = 4(180-x) Distribute 4 through, x = 720-4x Add 4x to both sides, 5x = 720, divide by 5, x = 144.

OpenStudy (anonymous):

==> 5x=720 ==> x=720/5 =144

OpenStudy (anonymous):

180-x = 4x, 5x = 180 -> x= 36

OpenStudy (anonymous):

It could be interpreted both ways... which angle is such that it's supplement is four times the angle itself. Does this mean the supplement is four times the angle? or is the angle four times the supplement?

OpenStudy (anonymous):

Determine which angle (x) is such that it's supplement (180-x) is four times the angle (x) itself.

OpenStudy (anonymous):

I'm thinking the answer would be 36, because it's saying that the supplement is 4x the other.

OpenStudy (anonymous):

Thanks, everyone, though.